# CHIMERA: A Knowledge Base of Idea Recombination in Scientific Literature

### v3

@@ Abstract
A hallmark of human innovation is the process of recombination—creating original ideas by integrating elements of existing mechanisms and concepts. In this work, we automatically mine the scientific literature and build CHIMERA: a large-scale knowledge base (KB) of recombination examples. CHIMERA can be used to empirically explore at scale how scientists recombine concepts and take inspiration from different areas, or to train supervised machine learning models that learn to predict new creative cross-domain directions. To build this KB, we present a novel information extraction task of extracting recombination from scientific paper abstracts, collect a high-quality corpus of hundreds of manually annotated abstracts, and use it to train an LLM-based extraction model. The model is applied to a large corpus of papers in the AI domain, yielding a KB of over 28K recombination examples. We analyze CHIMERA to explore the properties of recombination in different subareas of AI. Finally, we train a scientific hypothesis generation model using the KB, which predicts new recombination directions that real-world researchers find inspiring. Our data and code are available at https://github.com/noy-sternlicht/CHIMERA-KB
CHIMERA: A Knowledge Base of Idea Recombination in Scientific Literature
Figure 1: We automatically extract thousands of recombination examples describing how scientists connect ideas in novel ways.
@@ 1 Introduction
Figure 2: 1) We start by collecting human-annotated recombination examples and use them to finetune an LLM for information extraction. 2) Next, we apply the fine-tuned LLM on the arXiv corpus and build a knowledge base of recombination examples. 3) Given a context string and a query concerning the recombination of a certain graph node, our recombination model suggests directions based on knowledge learned from the KB.
@@
Recombination—creating original conceptual and physical combinations of existing mechanisms, methods, perspectives and approaches to address problems—is a common form of ideation Uzzi et al. (2013); Youn et al. (2015); Shi and Evans (2023). Recombination involves re-representing past ideas by decomposing them into conceptual chunks and then blending them into new solutions Knoblich et al. (1999); McCaffrey (2012), and also often involves forming abstract structural connections across domains Gentner et al. (1997); Gentner and Markman (1997); Gentner and Kurtz (2005); Chan et al. (2011); Frich et al. (2019)—for example, nature-inspired optimization algorithms that recombine concepts from nature and optimization.
In this work, we automatically mine CHIMERA, a knowledge base of recombination examples from across the scientific literature. We focus on two recombination types, which we name blends and inspirations. Blends combine multiple concepts to create new approaches (e.g., boosting classical machine learning algorithms using quantum computing), while inspirations involve adapting ideas from existing concepts to spark insight (e.g., applying bird flock behavior to coordinate drone swarms).
CHIMERA includes examples of blends of concepts within and across domains, and also inspirations in the form of analogies, reductions, and abstractions. Unlike simpler concept co-occurrence approaches Krenn et al. (2022) or more generic scientific extraction schema Luan et al. (2018), CHIMERA contains examples in which the authors explicitly mention a recombination as one of the core contributions of their work. Figure 1 presents a recombination example automatically extracted using our methods.
Our methods and collected data have broad potential applications in the field of Science of Science Shi and Evans (2019). The collected data enables empirical studies of innovation in novel ways—for example, examining how fields draw inspiration from one another and investigating how the blending of mechanisms across different domains emerges and evolves. Using our knowledge base, we present an analysis of sources of inspiration in the AI and NLP communities. We also demonstrate how to use this data to build a supervised learning framework for recombinant ideation. While existing work in the human-computer interaction (HCI) community provides researchers with tools for exploring idea recombinations Radensky et al. (2024a); Kang et al. (2022), CHIMERA allows us to take a different approach: training supervised models that learn from past examples how to recombine concepts for generating new scientific ideas. Figure 2 presents an overview of this process.
@@
Abstract: "…Current archaeology depends on trained experts to carry out bronze dating, which is time-consuming and labor-intensive. For such dating, in this study, we propose a learning-based approach to integrate advanced deep learning techniques and archaeological knowledge…"
@@
Abstract: "…Traditional approaches to enhance dialogue planning in LLMs, … either face efficiency issues or deliver suboptimal performance. Inspired by the dual-process theory in psychology, which identifies two distinct modes of thinking - intuitive (fast) and analytical (slow), we propose the Dual-Process Dialogue Planning (DPDP) framework…"
Inspiration: "the dual-process theory in psychology" $\longrightarrow$ "enhance dialogue planning in LLMs".
@@ 2 Related Work
Concept blending and finding inspiration through analogies are a key way to create new ideas McKeown (2014); 201 (2019); Holyoak and Thagard (1994). Recent research explores how idea recombination can enhance LLM-powered ideation tools. For example, CreativeConnect Choi et al. (2023) enables users to recombine keywords to generate graphic sketches, while Luminate Suh et al. (2023) facilitates the recombination of dimensional values to produce diverse LLM responses. Scideator Radensky et al. (2024b) is another recent work that allows researchers to explore new ideas by interactively recombining scientific concepts. Researchers have also investigated combining elements from input and analogous artifacts to create new ideas Srinivasan and Chan (2024); Chilton et al. (2019). In our work, we aim to extract a knowledge base of real recombinations in scientific papers, which can be used to facilitate research on recombination. For example, we train models that learn from past examples of recombination how to predict new recombinations (Figure 2).
@@
Much prior work has focused on information extraction (IE) from scientific papers. A notable example is the SciERC dataset Luan et al. (2018), which includes annotations of scientific entities (e.g., methods, tasks, metrics) and relations (e.g., conjunction) for 500 abstracts. Other, more recent examples are SciREX Jain et al. (2020), and SciDMTAL Pan et al. (2024), introducing document-level scientific IE datasets covering similar entity types. However, existing extraction approaches do not focus on recombination relationships, as we demonstrate in Appendix L, Figure 18. In our work we design a simple information extraction schema that is focused on idea recombination, to enable exploration of this important form of scientific innovation. For example, as part of our knowledge base we are able to extract many examples of analogical inspirations used by the AI community (Figure 1), which was not possible using existing scientific IE schema.
@@ 3 Extracting Recombinations
3.1 Problem Definition
We focus on cases where paper abstracts clearly say that a contributed idea is rooted in an original combination of concepts or explicitly mention a source of inspiration. As discussed in the introduction, we define two coarse-grained relation types to capture the notion of recombination: blend and inspiration. Blend refers to joining multiple concepts (e.g., methods, models, theories). Note that we use the terms concept blend and concept combination interchangeably. Inspiration refers to using knowledge from a source entity and implementing it in a different target entity. This could involve, for example, using an analogy or an abstraction as a source of inspiration, or more generally being influenced by another line of work. Relations are defined between free-form spans of text representing scientific concepts (see Figure 1, and additional examples in Table 1). We refer to the entities in a blend relation as combination-elements and those in an inspiration relation as inspiration-source and inspiration-target.
3.2 Recombination Mining
Our approach to mining recombination examples begins with building a curated dataset of annotated examples. We then use this dataset to train an information extraction model. Finally, we apply the trained model to collect recombination examples at scale. This process is illustrated in Figure 2.
Data sourcing
We use AI-related papers from the unarXive corpus Saier and Färber (2020) as a source of annotation examples [^1] We focus on the following arXiv categories: cs.AI, cs.CL, cs.CV, cs.CY, cs.HC, cs.IR, cs.LG, cs.RO, cs.SI. The data undergo an initial keyword-based filtering to identify works that are more likely to specify idea recombination. Table 7 in Appendix A lists the keywords used in this process. We then assign the remaining abstracts to annotators.
@@ 3.2 Recombination Mining
Table 2: Human-annotated corpus. We use examples with and without recombinations ("not-present"), simulating extraction over real-world data.
@@
After a thorough screening session, we recruit two annotators with scientific PhDs from Upwork [^2] https://www.upwork.com. The screening includes annotating examples following a detailed guidelines document [^3] https://tinyurl.com/zy27uhdp. Annotators who passed the screening had a personal one-hour training session to discuss additional examples. The annotators conduct the annotation through LightTag Perry (2021), a web-based text annotation platform (no longer operational). We collect a total of 580 annotated abstracts in this manner, as presented in Table 2.
To ensure annotation quality, we assign $10\%$ of the examples to both annotators and review the shared section at the end of each batch. During the review, we discuss disagreements and provide the annotators with feedback, after which they revise their work. The annotations undergo an additional review by an NLP expert. The expert verifies correctness, adjusts span boundaries, and merges annotations from different annotators.
Automatic recombination mining
We next use the collected data to fine-tune an LLM-based extraction model. Given the text and annotation schema, we instruct the model to extract the most salient recombination from the text, if one exists. The model must determine whether the text discusses recombination, infer its type, and identify entities in a single query.
We devise the test set from examples where at least two annotators (out of three) agree on the recombination type (or lack of recombination). This provides high-quality test data with fewer ambiguities. Table 2 presents the distribution of example types in both the train and test sets. Additional implementation details related to the extraction baselines are discussed in Appendix B.
@@
Table 3: The CHIMERA knowledge base summary. Our knowledge base encompasses over 28K recombination examples, a quarter of which are interdisciplinary.
@@ 3.3 The CHIMERA Knowledge Base
This section describes building the CHIMERA knowledge base. We first use our extraction method to mine recombination examples, categorize them, and build a KB in which nodes represent scientific concepts, and edges represent recombination relations between them.
@@
We source abstracts from the arXiv dataset [^4] https://tinyurl.com/mrzksbky. This dataset updates monthly and includes newer examples than unarXive Saier and Färber (2020). We apply our fine-tuned extraction model over publications from 2019 to 2024 within the same CS categories used for the annotation task. After applying the model we filter predictions that fail to comply with the data schema or could not be properly parsed.
Categorization
In addition to extracting the relations, we apply GPT-4o to identify the scientific domain of each extracted entity given the abstract. This enables analyses we perform in Section 4.2. The model is instructed to select the most appropriate arXiv category from either the arXiv taxonomy list or a supplementary list of non-arXiv domains (e.g., "psychology"). In cases where no given label captures the entity’s scientific taxonomy, we classify the node’s domain as Other. The analysis in Section 4.2 omits nodes from this domain, as we find they might be too noisy, too general or miscellaneous. Examples of such nodes, along with the used prompts and additional technical information about this step, are available in Appendix D.
We further assign the graph nodes a higher-level scientific discipline. The discipline is either the arXiv group name if available ("computer-science" for cs.AI), or the relevant non-arXiv domain.
@@
Does it discusses recombination?
Human-agreement 0.757 0.765 0.760
@@
We normalize the knowledge base entities by clustering semantically similar ones, and further enrich each edge in the graph with the publication date and arXiv categories of the paper citing it. For simplicity, we focus on binary relations when building the graph. Table 3 reports the summary of the final KB, including the number of interdisciplinary blends and inspirations.
(a) Frequent inspiration source and target domains.
(b) Frequent domains of blend nodes.
(c) Common sources of inspiration for leading domains.
Figure 3: Recombinations between areas. cs.*, q-bio.nc and math.oc are arXiv categories. Inspirational connections are often cross-domain (Figure 3(a)), while blends more often stay within-domain (Figure 3(b)). Figure 3(c) zooms in on a few domains, for example revealing that robotics often draws inspiration from zoology).
@@ 4 Results
4.1 Recombination extraction evaluation
We evaluate the recombination extraction process in three different levels of increasing difficulty: abstract classification (whether the text discusses recombination), entity extraction (what entities are described in the text) and relation extraction (what the discussed relation is). To evaluate abstract classification, we use precision, recall and F1. We use a soft evaluation metric for entity and relation extraction tasks, where two entities of the same type match if they refer to a semantically similar concept. We utilize GPT-4o-mini as a judge of content similarity (Figure 12 in Appendix E presents our prompt). We select GPT-4o-mini over GPT-4o after conducting a qualitative examination and finding only a handful of cases in which the model judgment differs (3 span pairs in the entire test set). To avoid position bias, we run the judge twice for each pair, reversing the span order each time. We consider two spans equivalent only when the judge returns a positive answer on both runs. Each predicted entity can match with exactly one gold entity and vice versa, with any additional matches being ignored. Under this definition of soft entity matching, we compute the entity extraction quality of a model using standard precision, recall and F1. For recombination relations, we apply the same metrics using partial relation matching: a predicted relation’s contribution to the true-positive score depends on how many entities match with a ground-truth relation of the same type.
Extraction results
Table 4 reports results for different levels of recombination extraction (abstract classification, entity extraction and relation extraction). We experiment with end-to-end (E2E) extraction approaches, inferring whether the text discusses recombination, its type, and what entities are involved - all at once. Note that when handling E2E approaches, we regard any extracted relation as a positive abstract classification. We also study models specialized in classification (Abstract-classifiers) or entity extraction (Entity-extractors). Appendix B describes implementation details regarding extraction.
Human agreement has F1 scores of $0.760$, $0.675$, and $0.651$ for classification, entity extraction, and relation extraction, respectively. These values are comparable with other works measuring soft annotators’ agreement in complex extraction tasks Naik et al. (2023); Sharif et al. (2024).

### v4

# CHIMERA: A Knowledge Base of Scientific Idea Recombinations for Research Analysis and Ideation

@@ Abstract
A hallmark of human innovation is recombination—the creation of novel ideas by integrating elements from existing concepts and mechanisms. In this work, we introduce CHIMERA, a large-scale Knowledge Base (KB) of over 28K recombination examples automatically mined from the scientific literature. CHIMERA enables large-scale empirical analysis of how scientists recombine concepts and draw inspiration from different areas, and enables training models that propose novel, cross-disciplinary research directions. To construct this KB, we define a new information extraction task: identifying recombination instances in scientific abstracts. We curate a high-quality, expert-annotated dataset and use it to fine-tune a large language model, which we apply to a broad corpus of AI papers. We showcase the utility of CHIMERA through two applications. First, we analyze patterns of recombination across AI subfields. Second, we train a scientific hypothesis generation model using the KB, showing that it can propose novel research directions that researchers rate as inspiring. We release our data and code at https://github.com/noy-sternlicht/CHIMERA-KB.
CHIMERA: A Knowledge Base of Scientific Idea Recombinations for Research Analysis and Ideation
Figure 1: We propose a new task of extracting recombinations: examples of how scientists connect ideas in novel ways. The extracted information enables applications in research analysis and automated ideation.
@@ 1 Introduction
Figure 2: CHIMERA KB construction and applications. Construction: (1) We use human-annotated recombination examples to fine-tune an LLM for information extraction; (2) the model extracts recombinations from arXiv abstracts to build a large-scale KB. Applications: CHIMERA supports diverse use cases, including computational ideation, exploration of recombination patterns across scientific domains, and meta-scientific analysis.
@@
Recombination—the creation of novel conceptual or physical solutions by combining existing mechanisms, methods, perspectives—is a widely recognized mechanism of ideation and innovation Uzzi et al. (2013); Youn et al. (2015); Shi and Evans (2023). It involves reinterpreting prior ideas by breaking them into components and blending them into new solutions Knoblich et al. (1999); McCaffrey (2012). This often requires forming abstract structural mappings across domains Gentner et al. (1997); Gentner and Markman (1997); Gentner and Kurtz (2005); Chan et al. (2011); Frich et al. (2019)—e.g., as in bio-inspired algorithms that apply biological principles to computational problems.
In this work, we introduce a new task: extracting recombinations from scientific papers. We present CHIMERA, a large-scale knowledge base (KB) of recombination examples automatically mined from papers. Figure 1 shows one such case, where a robotic design is inspired by animal mechanics. CHIMERA enables exploring, analyzing, and training models on such examples, capturing a fundamental pattern of human ingenuity.
Unlike simpler concept co-occurrence methods Krenn et al. (2022) or general scientific extraction schemas Luan et al. (2018), CHIMERA targets cases where authors explicitly describe recombination as central to their contribution. We focus on two broad recombination types: blends, which combine concepts into novel approaches (e.g., augmenting classical ML with quantum computing), and inspirations, where ideas from one domain spark solutions in another (e.g., using bird flock behavior to coordinate drones). CHIMERA captures both intra- and cross-domain cases, including analogies, abstractions, and reductions.
The resulting KB and methods enable diverse uses (Figure 2). In this paper, we focus on two applications that have seen growing interest in recent years: Science Analysis Fortunato et al. (2018); Wahle et al. (2023); Pramanick et al. (2025) and Scientific Ideation Wang et al. (2024); Si et al. (2024); Radensky et al. (2024); Garikaparthi et al. (2025).
Science Analysis. We demonstrate how CHIMERA supports meta-scientific analysis (also known as science of science or scientometrics) Fortunato et al. (2018): empirical studies of how innovation unfolds. Researchers conducting meta-science analyses aim to understand how a field (e.g., AI) evolves over time and identify trends (e.g., emerging connections across areas). CHIMERA allows analysis of how ideas are combined within and across domains Shi and Evans (2019), and of how disciplines, topics and concepts inspire one another. This provides a direct and precise alternative to traditional citation-based Wang et al. (2015); Myers et al. (2013); Wahle et al. (2023) or co-occurrence-based approaches Frohnert et al. (2024), which are often coarse and noisy. Unlike these methods, CHIMERA allows to identify how a scientific idea is formed by blending concepts or by taking inspiration from another concept, unlocking new and also more granular analyses.
Scientific Ideation. We show how CHIMERA supports training and evaluating scientific hypothesis generation models Wang et al. (2024), by learning from patterns of past recombinations to propose novel concept blends or inspirations (e.g., new analogical inspirations). Prior work has explored suggesting analogical recombinations via unsupervised discovery Radensky et al. (2024); Hope et al. (2017); in contrast, CHIMERA provides the first large-scale resource with real, author-described examples of how research problems were addressed via recombination. This enables supervised recombination models to observe many examples of how recombinations have been applied to specific problems (e.g., the cross-domain inspiration in Figure 1), and learn to suggest relevant blends or inspiration directions for new problems.
Finally, CHIMERA also enables faceted search and exploration Katz et al. (2024). Researchers can search the KB to find cases of cross-domain inspirations within a topic of interest (e.g., search for all robotics ideas inspired by zoology), sparking new creative directions.
To conclude, our contributions are as follows:
  • We present CHIMERA, the first knowledge base of idea recombination examples described by authors in scientific papers. CHIMERA distinguishes between two core types: blends and inspirations, enabling nuanced analysis in downstream tasks.
  • We define a novel extraction task to identify recombinations in scientific abstracts, and release a high-quality, expert-verified dataset of 500+500+500 + manually annotated examples, accompanied by fine-tuned extraction baselines.
  • We show CHIMERA’s utility through two applications: a) Meta-scientific analysis of recombination patterns, and b) Computational ideation, where models trained on CHIMERA propose novel recombination directions.
@@
Abstract: “…Current archaeology depends on trained experts to carry out bronze dating… we propose to integrate advanced deep learning techniques and archaeological knowledge…”
@@
Abstract: “…Traditional approaches to enhance dialogue planning in LLMs, … either face efficiency issues or deliver suboptimal performance. Inspired by the dual-process theory in psychology… we propose the Dual-Process Dialogue Planning (DPDP) framework…”
[…]
@@ 2 Related Work
Blending concepts and analogical inspiration are core mechanisms of ideation and innovation in cognitive science and creativity research McKeown (2014); 201 (2019); Holyoak and Thagard (1994). These processes involve combining or re-representing existing ideas to produce novel concepts and solutions.
Recent work explores how idea recombination can enhance LLM-powered ideation tools. For example, CreativeConnect Choi et al. (2023) lets users recombine keywords to generate graphic sketches, while Luminate Suh et al. (2023) supports recombination of dimensional values to produce diverse LLM responses. Scideator Radensky et al. (2024) is another recent work that helps researchers explore ideas through interactive concept recombination. Other studies focus on recombining ideas from input and analogous artifacts Srinivasan and Chan (2024); Chilton et al. (2019) or searching for useful recombinations via iterative idea generation Yang et al. (2025a, b).
In this work, we build CHIMERA, the first KB of scientific idea recombinations, and show how it enables a new approach for recombinant ideation: training models that learn from past examples of how ideas have been recombined in scientific texts, to suggest new recombination directions.
@@
Information extraction (IE) from scientific texts has been widely studied in NLP. A foundational resource is SciERC Luan et al. (2018), which labels scientific entities (e.g., methods, tasks, metrics) and generic relations (e.g., conjunction) across 500500500 abstracts. Later datasets, such as SciREX Jain et al. (2020) and SciDMTAL Pan et al. (2024), expand IE to full documents, but similarly focus on standard schema involving scientific concepts and their relations. However, existing extraction approaches are not designed to capture recombination relationships, often resulting in noisy, irrelevant, or misleading outputs, as we illustrate in Appendix G, Figure 20.
In this work, we introduce a focused IE schema tailored specifically to idea recombination, along with a taxonomy that distinguishes between key recombination types: blend and inspiration. This enables a more precise and semantically rich analysis of cross-domain ideation. For instance, our knowledge base includes numerous analogical inspirations identified in AI research (Figure 1) - patterns that existing scientific IE schemas fail to capture.
@@ 3 Extracting Recombinations
Problem definition
We focus on scientific abstracts where authors explicitly link their contribution to a novel combination or clear source of inspiration. As outlined in the introduction, we capture this with two coarse-grained relation types: blend and inspiration. Blend refers to the fusion of multiple concepts–such as methods, models, or theories–into a new solution or framework. We use the terms “concept blend” and “concept combination” interchangeably. Inspiration, by contrast, refers to transferring knowledge or insight from one entity (the source) to another (the target). This transfer may be realized through analogies, abstraction, or more general links to influential prior work.
Each relation is defined over free-form text spans that represent scientific concepts (see Figure 1; additional examples in Table 1). In blend relations, we refer to the participating entities as combination-elements; in inspiration relations, we refer to them as the inspiration-source and inspiration-target. This schema captures diverse recombination phenomena, such as metaphor, reduction, or abstraction (as illustrated in Appendix D.2) while remaining conceptually clear and efficient to annotate. It offers practical annotation advantages and strong alignment with ideation theory McKeown (2014); 201 (2019); Holyoak and Thagard (1994).
@@
Table 2: Human-annotated corpus. We include also negative examples without recombinations (“not-present”).
3.1 Recombination Mining
We begin by curating a dataset of annotated recombination examples, which we use to train an information extraction model. The trained model is then applied to extract recombinations at scale. This process is illustrated in Figure 2.
Data sourcing
We annotate AI-related papers from the unarXive corpus Saier and Färber (2020) [^1] We focus on the following arXiv categories: cs.AI, cs.CL, cs.CV, cs.CY, cs.HC, cs.IR, cs.LG, cs.RO, cs.SI. The data undergo an initial keyword-based filtering to identify works that are more likely to specify idea recombination. Table 8 in Appendix B.1 lists the keywords used in this process. We then assign the remaining abstracts to annotators.
@@ 3.1 Recombination Mining
Our annotation setup follows standard IE practices, using two trained annotators and expert review to balance quality and feasibility Naik et al. (2024); Sharif et al. (2024); Pramanick et al. (2025). Following a screening phase, we recruited two annotators with scientific PhDs via Upwork [^2] https://www.upwork.com, selected from a pool of highly experienced workers we had previously collaborated with. Screening involved annotating examples using a detailed guidelines document [^3] https://tinyurl.com/zy27uhdp, followed by a one-hour training session covering additional examples and edge cases. Annotation was conducted using LightTag Perry (2021), a web-based annotation platform. This process yielded 580580580 annotated abstracts, summarized in Table 2. To monitor annotation quality, we assign 10% of the examples to both annotators and review this shared subset after each batch. Disagreements are resolved through discussion and revision. All annotations are then reviewed by an NLP expert, who verifies correctness, refines spans, and consolidates annotations.
@@
Table 3: CHIMERA contains over 28K recombinations, a quarter of them interdisciplinary.
Automatic recombination mining
We use the collected data to fine-tune an LLM-based extraction model. We instruct the model to extract the most salient recombination from the text, if one exists. The model must determine whether the text discusses recombination, infer its type, and identify entities in a single query. We devise the test set from examples where at least two annotators (out of three) agree on the recombination type (or absence), ensuring high-quality, low-ambiguity data. Table 2 summarizes the train and test sets.
@@ 3.2 The CHIMERA Knowledge Base
We construct the CHIMERA knowledge base by mining recombination examples from scientific abstracts, categorizing them, and representing them in a graph where nodes are scientific concepts and edges denote recombination relations.
@@
We use abstracts from the arXiv dataset [^4] https://tinyurl.com/mrzksbky, which updates monthly and includes more recent papers than unarXive Saier and Färber (2020). We apply our fine-tuned extraction model over publications from 2019-2024 within the same CS categories used for the annotation task. We then filter out predictions that don’t conform to the data schema or cannot be parsed.
@@
Does it discuss a recombination?
Human-agreement 0.786 0.795 0.789
@@
Categorization
We apply GPT-4o to identify the scientific domain of each extracted entity given the abstract. This enables analyses we perform in Section 4.3. Further, each node is assigned a higher-level discipline—either the arXiv group name (e.g., “computer-science” for cs.AI) or a relevant non-arXiv domain. Additional technical details regarding this step appear in Appendix C.
@@
We normalize entities by clustering semantically similar ones. Next, we enrich each edge in the graph with the publication date and arXiv categories of the paper citing it. For simplicity, we focus on binary relations. Table 3 summarizes the resulting KB, including counts of interdisciplinary blends and inspirations.
@@ 4 Results
4.1 Experimental Settings
Evaluation criteria
We evaluate (1) Abstract classification–does the text discuss recombination?, (2) Entity extraction–what entities are described? and (3) Relation extraction–what is the relation discussed? For abstract classification, we report precision, recall, and F1. For entity and relation extraction, we adopt a soft matching approach: two entities of the same type match if they refer to semantically similar concepts. We use GPT-4o-mini [^5] Performed on par with GPT-4o. to judge similarity (see prompt and details in Appendix B.4).
A predicted entity may match at most one gold entity, and vice versa; extra matches are ignored. We compute precision, recall, and F1 under this soft matching. For relations, we use partial matching: a predicted relation contributes to the true positive count proportionally to the number of correctly matched entities in a gold relation of the same type.
We measure inter-annotator agreement using the same precision, recall, and F1 metrics, following standard practice in information extraction, where one annotator is treated as the gold reference Naik et al. (2023); Sharif et al. (2024).
(a) Frequent domains in inspiration edges.
(b) Frequent domains in blend edges.
(c) Common sources of inspiration in leading domains.
Figure 3: Recombinations between areas. cs.*, q-bio.nc and math.oc are arXiv categories. Inspirational connections are often cross-domain (Figure 3(a)), whereas blends tend to occur within the same domain (Figure 3(b)). Figure 3(c) zooms in on a few domains, for example, revealing that robotics often draws inspiration from zoology.
Extraction baselines
We evaluate several extraction baselines, including end-to-end (E2E) models that jointly predict whether an abstract discusses recombination, identify its type, and extract the involved entities. In these models, the prediction of any relation is treated as a positive signal for abstract-level classification. We also assess specialized models for individual sub-tasks: Abstract classifiers, which predict whether the text discusses recombination, and Entity extractors, identify relevant entities. Implementation details for the baselines are in Appendix B.2. To contextualize model performance, we compare results against inter-annotator agreement, used as a proxy for human-level performance. Appendix A presents additional details concerning agreement computation.
4.2 Extraction Results
Table 4 reports results for abstract classification, entity extraction, and relation extraction. Human agreement scores are 0.7600.7600.760, 0.6750.6750.675, and 0.6510.6510.651 respectively, aligning with soft annotator agreement reported in similar complex extraction tasks Naik et al. (2023); Sharif et al. (2024). Cohen’s $\kappa$ also indicates moderate to substantial agreement: $\kappa=0.578$ for abstract classification, 0.6310.6310.631 for entity extraction, and 0.5420.5420.542 for relation extraction. Analysis of annotator disagreement is provided in Appendix A.1—most disagreements concern the presence of a recombination or the identification of its constituent entities, while disagreements over the recombination type are relatively rare.